OpenStudy (anonymous):

How does the strength of an electromagnet depend on the current and the number of turns in the coil?

OpenStudy (anonymous):

During a race, a runner runs at a speed of 6 m/s. 2 seconds later, she is running at a speed of 10 m/s. What is the runner’s acceleration?

OpenStudy (anonymous):

If you ride your bike at an average speed of 4 km/h and need to travel a total distance of 28 km, how long will it take you to reach your destination?

OpenStudy (anonymous):

A small dog is trained to jump straight up a distance of 1.2 m. How much kinetic energy does the 7.2-kg dog need to jump this high? (The acceleration due to gravity is 9.8 m/s2.) Show your work.

OpenStudy (anonymous):

As 390 g of hot milk cools in a mug, it transfers 30,000 J of heat to the environment. What is the temperature change of the milk? The specific heat of milk is 3.9 J/g•°C. Show your work.

OpenStudy (anonymous):

1. A solenoid is a electromagnet. From \[B=\mu \times n \times i\] B-strength of magnetic field. u-permeability of the core. n-number of turns. i-current. You can see why 'B' depends on 'i' and 'n'.

OpenStudy (anonymous):

2. If the acceleration is constant , a=10-6/2 i.e \[2 m/s ^{2}\]

OpenStudy (anonymous):

3. The definition of avg.speed is total distance traveled divided by the total time taken.

OpenStudy (anonymous):

4. The kinetic energy is getting converted to potential energy. So, decrease in kinetic energy=increase in potential energy.\[1/2\times m \times v ^{2} = m \times g \times h\] Now mgh is your answer.

OpenStudy (anonymous):

\[Q=m \times c \times \Delta T\] Q- heat exchanged. m- mass. c- specific heat. delta(T)-change in temperature.